### v2

@@ 4 The BeCoS Benchmark
Datasets. BeCoS is composed of seven different datasets with four-legged creatures and humanoid shapes encompassing challenging and diverse shape matching scenarios for both isometric and non-isometric deformations: TOSCA [bronstein2008numerical], FAUST [bogo2014faust], SCAPE [anguelov2005scape], KIDS [rodola2014dense], DT4D [magnet2022smooth], SMAL [Zuffi:CVPR:2017] and SHREC’20 [dyke2020shrec]. See Table 1 for more information about the used datasets. From these datasets, we use a total of 2543 shapes (496 are humanoid shapes and 2047 are four-legged shapes) to generate pairwise matching problems. We partition our dataset into predefined training/validation/test splits, ensuring evaluation is performed on unseen categories. To combat dataset imbalance, we oversample the smaller datasets in our training set. The number of shape matching instances in the train/validation/test splits is 10185/137/142. We limit the test set size to easily compare to axiomatic methods, which are often computationally heavy.
Predefined Benchmark Parameters. We choose the following parameters in the data generation pipeline to create a challenging but reasonable benchmark. For re-meshing, we use quadric decimation [garland1997surface] and we fix the range of vertices to a random number between $9000{-}10{,}000$ to facilitate training of current supervised/unsupervised methods and limit memory consumption. Note that re-meshing the same shape multiple times results in different re-meshed versions of the shape due to the random selection of the number of vertices. Our main setting introduces a random rotation around the z-axis, assuming each shape has a natural ‘upright’ direction. The maximum camera angle for the partial-to-partial setting is set to $\alpha=\frac{\pi}{4}$. Similar to [attaiki2021dpfm] and [ehm2024partial], these parameters are chosen to maintain an overlap of $10\%-90\%$, using $m=10$ maximum iterations (see supp. mat. Sec. 9.3 for distribution of overlap). All default parameters to generate the BeCoS benchmark are explicitly stated and explained in details in our implementation repository.
@@
GeomFMaps [donati2020deepGeoMaps] Sup wks, xyz
DPFM [attaiki2021dpfm] Sup/Unsup wks, xyz
ULRSSM [cao2023unsupervised] Unsup wks, xyz
Smooth Shells [eisenberger2020] Axio shot
PFM [rodola2017partial] Axio shot
SM-COMB [roetzer2022scalable] Axio xyz
GC-PPSM  [ehm2024partial] Axio xyz
@@
Full-to-Full (F2F) Source GT GeomFMaps ULRSSM Smooth Shells
@@
Partial-to-Full (P2F) Source GT DPFM (Unsup) ULRSSM PFM
@@
Partial-to-Partial (P2P) Source GT DPFM (Sup) SM-COMB GC-PPSM
@@
Figure 4: Cross-dataset and cross-category shape matching results visualised via colour transfer (red denotes missing parts). All baselines show sub-optimal performance on all three settings (some examples are highlighted with blue circles). The green square shows the semantic missing parts between the source and target mesh. As GC-PPSM and SM-COMB are solved on very low resolution when upsampling, we do nearest neighbour search in the features space of the matched patches (see [ehm2024partial] for more information), ensuring geometric consistency in low resolution. Both methods show patch-wise inaccuracies due to this low resolution solution.
@@ 5.1 Evaluation Metrics
Geodesic Error. We evaluate the geodesic error of correspondences using the Princeton Protocol [kim2011blended]. This evaluation entails normalising the geodesic error between the ground truth and the computed correspondences by the shape diameter, defined as the square root of the area of the target shape. For partial shapes, we use the square root of the area of their corresponding full shape. To handle the partial-to-partial matching paradigm, we follow [ehm2024partial] and set an infinite geodesic error for unmatched vertices that should be matched and for vertices that are matched but should not be. In contrast to the partial-to-full (P2F) and full-to-full (F2F) settings, where no overlapping region is determined, the overlapping region prediction affects the geodesic error curve for the partial-to-partial setting.
@@
where $\chi_{i}$ is the left/right prediction for the $i$-th vertex in $\mathcal{X}$, $\mathds{1}$ is the indicator function, $\chi_{i}^{gt}$ is the ground truth of the left/right annotation of the $i$-th vertex in $\mathcal{X}$. We use a similar approach for the partial-to-partial case as when using the geodesic error. For unmatched vertices that should be matched and vertices that are matched but should not be, we set an accuracy of 00.
@@ 5.4 Ablation Studies
We select the partial-to-full methods for this ablation, as the partial-to-partial methods perform prohibitively poorly, resulting in a mostly random prediction accuracy. Many full-to-full methods use wave kernel signature features [aubry_wave_2011] as input, making them independent of rotations.
@@ 6 Discussion and Future Work
The cross-category and cross-dataset correspondences of BeCoS are generated pairwise for the selected template shapes. Although they are of good quality, for cross-category shapes they may give rise to correspondence ambiguity (e.g. hooves to paws). Ambiguities may potentially accumulate if many pairwise ground truth correspondences are chained (i.e. for shapes that are distant in our shape network of shapes), which deteriorates the composed ground truth quality (see left example in Figure 8 in the supplementary material). This could potentially be improved by incorporating cycle-consistency or synchronisation techniques [huang2013consistent, pachauri2013solving, bernard2021sparse] in the process of generating ground truth correspondences.
@@ Acknowledgments
[…]
Beyond Complete Shapes: A Benchmark for Quantitative Evaluation
of 3D Shape Surface Matching Algorithms
@@ 8.1 Alternative Energies on the Partial-to-Partial Setting
In the main paper, we use XYZ coordinates for the axiomatic methods SM-COMB [roetzer2022scalable] and GC-PPSM [ehm2024partial] for the partial-to-partial setting. While these methods work best with learned features (e.g. unsupervised features from ULRSSM [cao2023unsupervised]), the latter features are not reliable for our challenging benchmark dataset (see Fig. 5 in the main paper for the performance of ULRSSM in the partial-to-full setting). In our case, we found that XYZ coordinates as feature descriptors are most suited for our benchmark. We compare the XYZ features to the original energy formulation used for SM-COMB (a combination of bending and membrane energy) and the rotation-invariant SHOT features. We show the results in Fig. 7 for the geodesic distance comparison and Table 6 for the overlapping region prediction. Both show that XYZ as input features work best in the given setting. This is the case because in BeCoS the shapes are rotated only around the z-axis, which leaves the z-coordinates as reliable information. While this is not always a realistic setting, as it would not work for completely random rotations, we choose this setting to be able to evaluate the axiomatic methods in a (reasonably) meaningful manner. As these methods will benefit from better features, we hope to encourage further research in this direction.
@@ 9.1 Correspondence Generation
This procedure is similar to the one used by [magnet2022smooth], and we use the FaceForm software with an academic license [faceform]. The steps of our procedure can be summarised as follows:
@@
    To get the correspondences between the two shapes, we find for every vertex $\hat{A}$ the closest triangle of template shape $B$. We then project every vertex of $\hat{A}$ on the closest triangle and return the barycentric coordinates. The correspondences of semantic missing parts are set to $-1$.
@@ 10 Details about the BeCoS Benchmark and the Framework
As mentioned in the main paper, we use seven different full datasets in our benchmark. We use each dataset entirely with two exceptions: For TOSCA we exclude the gorilla as it is excluded from most shape matching analyses due to its difficulty. The original dataset [li20214dcomplete] of DT4D provides animated meshes for each category with a lot of categories. Some categories in the original dataset are difficult to semantically match to other categories, e.g. dragons. Therefore, we use the same categories that were selected from DT4D-Matching [magnet2022smooth]. This dataset provides its own cross-dataset correspondences but only for a version of the dataset where each mesh is around 8000 vertices. Since the shapes are only re-meshed and not in the original resolution, we use the original animated files but extract the same categories and frames from the animation as DT4D-Matching. This ensures that we have higher resolution input shapes and less error propagation along the correspondence paths, as the shapes in the same category have the same triangulation. We end up with 39 categories for four-legged animals and eight categories of humanoid shapes.
@@ 12 Compute and Resources
DPFM [attaiki2021dpfm] 8-20h 1.1s
SM-COMB [roetzer2022scalable] N/A 0.08h
GC-PPSM [ehm2024partial] N/A 2.5h
PFM [rodola2017partial] N/A 0.2h
ULRSSM [cao2023unsupervised] 16-86h 20s
Smooth Shells [eisenberger2020] N/A 0.2h
GeomFMaps [donati2020deepGeoMaps] 3-7h 1.1s

### v3

@@ 4 The BeCoS Benchmark
Datasets. BeCoS is composed of seven different datasets with four-legged creatures and humanoid shapes encompassing challenging and diverse shape matching scenarios for both isometric and non-isometric deformations: TOSCA [BBK08], FAUST [BRLB14], SCAPE [ASK*05], KIDS [RRW*14], DT4D [MRSO22], SMAL [ZKJB17] and SHREC’20 [DLR*20]. See Table 1 for more information about the used datasets. From these datasets, we use a total of 2543 shapes (496 are humanoid shapes and 2047 are four-legged shapes) to generate pairwise matching problems. We partition our dataset into predefined training/validation/test splits, ensuring evaluation is performed on unseen categories. To combat dataset imbalance, we oversample the smaller datasets in our training set. The number of shape matching instances in the train/validation/test splits is 10185/137/142. We limit the test set size to easily compare to axiomatic methods, which are often computationally heavy.
Predefined Benchmark Parameters. We choose the following parameters in the data generation pipeline to create a challenging but reasonable benchmark. For re-meshing, we use quadric decimation [GH97] and we fix the range of vertices to a random number between 900010,0009000{-}10{,}000 to facilitate training of current supervised/unsupervised methods and limit memory consumption. Note that re-meshing the same shape multiple times results in different re-meshed versions of the shape due to the random selection of the number of vertices. Our main setting introduces a random rotation around the z-axis, assuming each shape has a natural ‘upright’ direction. The maximum camera angle for the partial-to-partial setting is set to $\alpha=\frac{\pi}{4}$. Similar to [APO21] and [EGR*24], these parameters are chosen to maintain an overlap of $10\%-90\%$, using $m=10$ maximum iterations (see supp. mat. Sec. 9.3 for distribution of overlap). All default parameters to generate the BeCoS benchmark are explicitly stated and explained in details in our implementation repository.
@@
GeomFMaps [DSO20] Sup wks, xyz
DPFM [APO21] Sup/Unsup wks, xyz
ULRSSM [CRB23] Unsup wks, xyz
Smooth Shells [ELC20] Axio shot
PFM [RCB*17] Axio shot
SM-COMB [RSCB22] Axio xyz
GC-PPSM  [EGR*24] Axio xyz
@@
Full-to-Full (F2F)
Source GT GeomFMaps ULRSSM Smooth Shells
@@
Partial-to-Full (P2F)
Source GT DPFM (Unsup) ULRSSM PFM
@@
Partial-to-Partial (P2P)
Source GT DPFM (Sup) SM-COMB GC-PPSM
@@
Figure 4: Cross-dataset and cross-category shape matching results visualised via colour transfer (red denotes missing parts). All baselines show sub-optimal performance on all three settings (some examples are highlighted with blue circles). The green square shows the semantic missing parts between the source and target mesh. As GC-PPSM and SM-COMB are solved on very low resolution when upsampling, we do nearest neighbour search in the features space of the matched patches (see [EGR*24] for more information), ensuring geometric consistency in low resolution. Both methods show patch-wise inaccuracies due to this low resolution solution.
@@ 5.1 Evaluation Metrics
Geodesic Error. We evaluate the geodesic error of correspondences using the Princeton Protocol [KLF11]. This evaluation entails normalising the geodesic error between the ground truth and the computed correspondences by the shape diameter, defined as the square root of the area of the target shape. For partial shapes, we use the square root of the area of their corresponding full shape. To handle the partial-to-partial matching paradigm, we follow [EGR*24] and set an infinite geodesic error for unmatched vertices that should be matched and for vertices that are matched but should not be. In contrast to the partial-to-full (P2F) and full-to-full (F2F) settings, where no overlapping region is determined, the overlapping region prediction affects the geodesic error curve for the partial-to-partial setting.
@@
where $\chi_{i}$ is the left/right prediction for the $i$-th vertex in $\mathcal{X}$, $\mathds{1}$ is the indicator function, $\chi_{i}^{gt}$ is the ground truth of the left/right annotation of the $i$-th vertex in $\mathcal{X}$. We use a similar approach for the partial-to-partial case as when using the geodesic error. For unmatched vertices that should be matched and vertices that are matched but should not be, we set an accuracy of 0.
@@ 5.4 Ablation Studies
We select the partial-to-full methods for this ablation, as the partial-to-partial methods perform prohibitively poorly, resulting in a mostly random prediction accuracy. Many full-to-full methods use wave kernel signature features [ASC11] as input, making them independent of rotations.
@@ 6 Discussion and Future Work
The cross-category and cross-dataset correspondences of BeCoS are generated pairwise for the selected template shapes. Although they are of good quality, for cross-category shapes they may give rise to correspondence ambiguity (e.g. hooves to paws). Ambiguities may potentially accumulate if many pairwise ground truth correspondences are chained (i.e. for shapes that are distant in our shape network of shapes), which deteriorates the composed ground truth quality (see left example in Figure 8 in the supplementary material). This could potentially be improved by incorporating cycle-consistency or synchronisation techniques [HG13, PKS13, BCT21] in the process of generating ground truth correspondences.
@@ 8.1 Alternative Energies on the Partial-to-Partial Setting
In the main paper, we use XYZ coordinates for the axiomatic methods SM-COMB [RSCB22] and GC-PPSM [EGR*24] for the partial-to-partial setting. While these methods work best with learned features (e.g. unsupervised features from ULRSSM [CRB23]), the latter features are not reliable for our challenging benchmark dataset (see Fig. 5 in the main paper for the performance of ULRSSM in the partial-to-full setting). In our case, we found that XYZ coordinates as feature descriptors are most suited for our benchmark. We compare the XYZ features to the original energy formulation used for SM-COMB (a combination of bending and membrane energy) and the rotation-invariant SHOT features. We show the results in Fig. 7 for the geodesic distance comparison and Table 6 for the overlapping region prediction. Both show that XYZ as input features work best in the given setting. This is the case because in BeCoS the shapes are rotated only around the z-axis, which leaves the z-coordinates as reliable information. While this is not always a realistic setting, as it would not work for completely random rotations, we choose this setting to be able to evaluate the axiomatic methods in a (reasonably) meaningful manner. As these methods will benefit from better features, we hope to encourage further research in this direction.
@@ 9.1 Correspondence Generation
This procedure is similar to the one used by [MRSO22], and we use the FaceForm software with an academic license [Fac23]. The steps of our procedure can be summarised as follows:
@@
    To get the correspondences between the two shapes, we find for every vertex $\hat{A}$ the closest triangle of template shape $B$. We then project every vertex of $\hat{A}$ on the closest triangle and return the barycentric coordinates. The correspondences of semantic missing parts are set to 1-1.
@@ 10 Details about the BeCoS Benchmark and the Framework
As mentioned in the main paper, we use seven different full datasets in our benchmark. We use each dataset entirely with two exceptions: For TOSCA we exclude the gorilla as it is excluded from most shape matching analyses due to its difficulty. The original dataset [LTT*21] of DT4D provides animated meshes for each category with a lot of categories. Some categories in the original dataset are difficult to semantically match to other categories, e.g. dragons. Therefore, we use the same categories that were selected from DT4D-Matching [MRSO22]. This dataset provides its own cross-dataset correspondences but only for a version of the dataset where each mesh is around 8000 vertices. Since the shapes are only re-meshed and not in the original resolution, we use the original animated files but extract the same categories and frames from the animation as DT4D-Matching. This ensures that we have higher resolution input shapes and less error propagation along the correspondence paths, as the shapes in the same category have the same triangulation. We end up with 39 categories for four-legged animals and eight categories of humanoid shapes.
@@ 12 Compute and Resources
DPFM [APO21] 8-20h 1.1s
SM-COMB [RSCB22] N/A 0.08h
GC-PPSM [EGR*24] N/A 2.5h
PFM [RCB*17] N/A 0.2h
ULRSSM [CRB23] 16-86h 20s
Smooth Shells [ELC20] N/A 0.2h
GeomFMaps [DSO20] 3-7h 1.1s